OpenStudy (anonymous):

Find the perimeter of a rectangle if the length is 4mm and the width is 7mm show all you work.

OpenStudy (anonymous):

so do you know the formula for perimeter..?

OpenStudy (anonymous):

P=2l+2w l is the length and w is the width

OpenStudy (anonymous):

so that is all it was that easy

OpenStudy (anonymous):

yes so its P=2(4)+2(7)

OpenStudy (anonymous):

P=8+14 P=22mm

OpenStudy (anonymous):

what there are three what is the one???

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

is the answer is P=8+14 P=22mm

OpenStudy (anonymous):

yes but if you want to show all of you work you do P=2l+2w P=2(4)+2(7) P=8+14 P=22mm

OpenStudy (anonymous):

i need help with one more please

OpenStudy (anonymous):

Find the height of a triangle if the area is 30mm and the base is 12mm

OpenStudy (anonymous):

so do you know the formula for area of a triangle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[A=bh/2\]

OpenStudy (anonymous):

so we know A, which is area, and b, which is base

OpenStudy (anonymous):

so we plug those in, ill draw it one sec

OpenStudy (anonymous):